# Goldilocks Isolation: High Performance VMs with Edera

### v1

@@ 2.5 Kubernetes
Kubernetes [31] is a system for automatically deploying, scaling, and managing containerized applications. It runs bundled applications, known as container images, in pods that run on nodes. In the OS virtualization model above, a pod is equivalent to the application and nodes are equivalent to the virtual or physical machine running the shared kernel. A container runtime manages the lifecycle of pods. Kubernetes can work with any container runtime that follows the Container Runtime Interface (CRI) specification [30]. One such container runtime is runc, a lightweight OS virtualized runtime based on Docker. The Kubernetes control plane is responsible for deploying and managing pods. This is done primarily by the Kubernetes API server, which lets users query and change the state of Kubernetes. A kubelet is an agent that runs on each node that registers the node with the API server, and ensures that all pods expected by the API server are running.
@@
The Kubernetes ecosystem is composed of projects that add functionality, interacting with each other through defined APIs. These include container runtimes, image distribution services, and specialized control planes. The scope of this ecosystem means that any incompatible new technology will be at a huge disadvantage as it would lack the pre-built container images and services that can work with Kubernetes.
@@
Several systems have been built to allow Kubernetes to manage VMs instead of or in addition to OS virtualized containers. This provides the usability benefits of Kubernetes to hypervisor virtualization. Many of these, including Infranetes [28], rkt [38], and virtlet [41] require virtualization extensions. RunX [39] uses a series of shell scripts to run VMs on Xen, but tooling limitations make this project incompatible with current versions of Kubernetes. These projects demonstrate the desire for hypervisor virtualization systems compatible with Kubernetes, but none meet our performance and usability requirements.
2.6 GPU Driver Isolation
GPUs are increasingly used for general-purpose computation like machine learning and high performance computing, leading to an increased demand for GPUs in multi-tenant cloud environments. Multi-tenant environments reduce the cost of using a GPU as a single GPU may be shared among many users or applications. However, as with applications running on shared infrastructure, the shared GPU presents a new attack surface that includes GPU hardware and firmware. This GPU driver firmware is a complex piece of software with many known weaknesses [57, 13] despite efforts to improve their usability and security. Amazon EC2 [2] offers GPU instances that directly pass-through access to GPUs. However, each GPU is limited to a single VM. Some systems [14, 53] use a hypervisor to allow multiple VMs to access the same GPU. In these systems, the GPU device drivers run directly on the hypervisor, increasing the TCB. To achieve stronger isolation, some systems [17, 49] propose running GPU drivers within a trusted execution environment (TEE). Similarly, Graviton [55] proposes a design for a GPU TEE which requires hardware changes to the GPU, but provides full isolation of the GPU from the OS and CPU. None of these systems ensures that compromises in the GPU are fully isolated from the hypervisor without specialized hardware or changes to GPUs.
3 Requirements
Requirement Docker Unikernel gVisor Firecracker Kata Edera Edera hwv
Eliminate shared kernel $\times$× $\times$×
Existing images $\times$×
Runs on any commodity hw $\times$× $\times$× $\times$×
Driver isolation $\times$× $\times$× $\times$× $\times$× $\times$×
Memory safety $\times$× $\times$× $\times$× $\times$×
Runtime performance $\times$× $\times$×
Startup performance $\times$× $\times$× $\times$×
Table 1: We assess existing systems against our requirements. Docker using runc is included as a baseline as it uses OS virtualization. Edera hwv shows the results for Edera running on a system with hardware virtualization extensions.
The goal of Edera is to create strong container isolation, while ensuring usability. Usability means both the ability to take advantage of the large ecosystem built around containers, and the ability to adopt Edera without affecting runtime performance. To achieve this, our requirements are as follows.
@@ 3 Requirements
Eliminate the shared kernel. The system must support secure multi-tenancy, with the container acting as a security boundary. This means removing the kernel from the TCB so that a bug in the kernel is not sufficient to escape the isolation.
@@
Use existing images. Existing container images must run without modification. Some images have proprietary code which cannot be accessed by users to re-build the image. Even when the source is available, re-building images is time consuming.
Run on any commodity hardware. The system must be able to run on common, commercially available hardware without relying on specialized hardware like virtualization extensions or TEEs. Virtualization extensions are not available on all CPUs or on all commercial cloud infrastructure, and can be much more expensive when available. For example only 7% of AWS instance types include virtualization extensions.
Driver isolation. The system should be able to isolate hardware drivers, including those for GPUs, from the host kernel. In other words, a driver compromise should not be able to impact the hypervisor or host machine.
@@
Memory Safety. The system should be written in a memory safe language to reduce the attack surface of the hypervisor. Memory safety vastly reduces the instance of memory leaks and memory corruption attacks.
Runtime performance. The system’s runtime performance must be within 15% of running runc on Docker. This ensures that the performance is similar to that of systems using OS virtualization.
Startup performance. The system’s startup performance must be within 10% of that of runc on Docker. While the system startup occurs less often than runtime (just once per application), it is still important to have a fast startup time for provisioning new containers and scaling the number of copies up and down.
We analyze several systems discussed in section 2, and show how they compare to these requirements in table 1.
3.1 Threat model
@@
In multi-tenant environments, arbitrary workloads from third parties are run on shared infrastructure. The third party may be a company running their infrastructure for several customers, or a customer running their own containers. The infrastructure allows customers to run their own code (often in the form of Kubernetes containers), but the operator does not want customers to gain access to the underlying infrastructure or other customer workloads. Given this environment, for our threat model we assume an attacker can:
@@
  • Side channel attacks, for instance those that use the temperature of the host machine. These attacks are unavoidable on shared infrastructure, and are present in any virtual machine environment.
4 Design
@@
Edera is a type-1 paravirtualized hypervisor that runs workloads in zones, which are fully functional VMs that run standard Linux for container workloads. Edera has a root hardened zone that manages the hypervisor, passing along memory and CPU from the host machine. This root hardened zone can monitor for anomalies and shut down other zones if needed. The microkernel hypervisor controls the memory, page table, and CPU usage of zones. Each zone is scheduled like a user-space process within the hypervisor. Figure 2 provides an overview of Edera.
Figure 2: An overview of how Edera isolates zones.
4.1 Hypervisor
The Edera hypervisor is based on Xen’s hypervisor design [4], with a few key differences and additional features. For brevity this section focuses on the novel features of the Edera hypervisor. The Edera hypervisor is a microkernel hypervisor that acts as the host kernel and can run directly on the host machine. By using a microkernel, Edera minimizes the attack surface of the hypervisor. Research has shown that microkernels have fewer bugs, as many bugs come from less used pieces of the kernel [21].
Edera works in paravirtualization (PV) mode, with the hypervisor managing the CPU, MMU hardware page table, and network. There is also support for paravirualization with hardware virtualization (PVH) mode, which takes advantage of virtualization hardware for performance improvements. However, as PVH mode requires virtualization hardware, examples in this paper use PV mode unless otherwise stated. The page table mapping used by Edera is the same as in Xen, with the hypervisor giving guest OSs read-only access to virtual page tables that the hypervisor maps to the MMU hardware page table. CPU scheduling in the Edera hypervisor is more dynamic than the scheduling supported by Xen, with the resource manager able to dynamically pin and unpin CPUs for each zone. This dynamic configuration allows for more efficient resource allocation on the host machine as zones can be assigned CPUs and memory pages as needed, rather than these resources remaining static after zone creation.
4.2 Guest kernel
@@
Each zone launched by the hypervisor runs a guest OS kernel. Using PV, the hypervisor runs in ring 0, and all guest OSs run in a less privileged ring (ring 1 for x86 and ring 3 on x86-64) [44]. As most OSs expect to run in ring 0, the guest OS must be aware that it is running in a virtualized environment and support hypercalls. These hypercalls [4] are used for privileged operations that need to run in ring 0 and pass along those privileged system calls to the hypervisor. This is supported by many OSs through a PV mode.
Edera supports any Guest OS that has a PV mode, including most Linux distributions. For ease of use with Kubernetes, we create a custom guest OS kernel that is distributed as an OCI image. Thus, a user can specify this image in their Kubernetes configuration and automatically run their existing containers.
4.3 Inter-domain messaging
@@
Edera uses inter-domain messaging (IDM) to communicate with processes running in zones. IDM uses a Xen byte channel [42] to send protobuf messages to other zones using the transport layer. Protobuf (shortened from protocol buffers) is a platform-neutral mechanism for serializing structured data. Its design includes protection against overloading the pipe by using bound checks and other limits to ensure safe operation.
IDM is used to send commands or processes to the zone to be run (such as through the Kubernetes utility kubectl exec),in addition to sending logs and metadata back to the hypervisor. Any logs received from the IDM are considered untrusted as they come from a potentially-compromised user zone. However, they can be used to detect misbehavior in the zone.
4.4 Monitoring
@@
An init process in every zone starts when the zone is provisioned and runs a background process that monitors the VM. The init process sends all monitored events over the IDM to be logged and collected. If the init process is disabled, the hypervisor would mark the zone as not responding and deprovision it. As the init process runs on the zone, any compromise of the init process itself would only impact that zone.
@@
5 Implementation
Figure 3: An overview of the services used by Edera.
@@ 5 Implementation
We implemented Edera in Rust and MISRA C using the design from section 4 for both Kubernetes compatibility and GPU isolation. By implementing it in Rust and MISRA C, we avoid many memory safety bugs while achieving high performance. The implementation includes the hypervisor, root hardened zone, and several services that allow use with Kubernetes for cluster orchestration. These services include a Container Runtime Interface (CRI) compatible with the OCI runtime specification, daemons for zone monitoring, and a kube-apiserver for managing pods.
To start, we create a new set of Xen tooling in Rust to replace libxl. This tooling simplifies starting a new Xen domain by reducing the number of configuration files and introducing a Rust API to replace the command line tools used by libxl. On top of this new Xen tooling, we introduce several services.
5.1 Kubernetes
@@
We will walk through the system startup and running of a single Kubernetes pod in an Edera zone to illustrate the services used by the implementation, shown in fig. 3. First, the machine starts the microkernel when it is booted. This microkernel starts the hypervisor, which also acts as a Kubernetes host. The hypervisor launches protect-orchestrator, a management component that ensures that all other daemons are running. These daemons are protect-network, protect-daemon, and protect-cri.
@@
protect-daemon acts as the virtual machine monitor for the zones, keeping track of which zones are running each workload and managing drivers for the zones, while leaving resource management to the hypervisor. This is the daemon that creates, modifies, and removes zones. To do so, it manages an embedded key value store that tracks what is running in each zone. In this store each zone is identified by a uuid that is assigned on zone creation. In addition, protect-daemon manages the drivers for each zone, including the IDM. Using IDM, protect-daemon can pass commands from the API server (though protect-cri) to the init process in the zone. It can also receive messages from the zone, and pass these along to protect-cri.
@@
protect-orchestrator performs the initial system setup, and maintains running processes. It provides updates and health checks for running daemons including protect-daemon, protect-network, and protect-cri. In addition, it provides observability for all services running in the hypervisor, through tools like Prometheus [3].
@@
protect-network performs all packet routing and monitors network activity. Zones have a basic ethernet interface, but cannot perform packet routing without going through protect-network. protect-network receives tasks from zones through protect-daemon, then performs them on behalf of the zone. It writes packets to the networking layer and can dial sockets for zones. Zones do not have direct network access, so no packet gets routed without being seen by protect-network. This ensures that Edera can monitor all network activity.
@@
protect-cri provides a layer on top of the Kubernetes CRI, which is the protocol for communication between a kubelet and the container runtime. It can be used with any OCI-compliant CRI, and converts requests from the CRI to daemon requests for protect-daemon. protect-cri reconciles the state between protect-daemon and the CRI to ensure that the zones running on protect-daemon match those expected by the CRI. It also transfers messages from zones through protect-daemon to other Kubernetes processes. protect-cri supports running some Kubernetes nodes in Edera protect zones and others on different infrastructure. To do so, the protect-cri coordinates with containerd to run the non-Edera nodes.
@@
Once the system is booted, a user may launch a pod from protect-cri. This triggers the Kubernetes kubelet [18], which registers the pod with the kube-apiserver and launches a zone.
@@
Summary. In summary, Edera provides several services: protect-daemon, protect-orchestrator, protect-network, and protect-cri to manage zones and facilitate communication with Kubernetes. By routing all network traffic through protect-network and all IDM communication through protect-daemon, we ensure that all traffic can be monitored.
Figure 4: An overview of the use of Edera for GPU isolation. GPU drivers are moved from the hypervisor into a zone so that any misbehavior cannot effect the hypervisor or other zones.
5.2 Driver isolation
In addition to using Edera to interoperate with Kubernetes, we build device driver isolation using Edera zones. We demonstrate that Edera can be used for strong driver isolation that prevents device driver bugs from impacting the hypervisor. This driver isolation can be used along with the Kubernetes compatibility described above.
To illustrate driver isolation, we discuss the isolation of GPU drivers. Usually, drivers for GPUs and other hardware devices are run on the host kernel directly, then a sub-driver is exposed to each VM. If a VM requests an operation of the device, this operation is passed through to the driver on the host kernel. However, with this design, any bug in the device driver can lead to a corruption of the host kernel. Device drivers can be many gigabytes of proprietary code and have had bugs discovered in the past, with more than 20 CVEs in 2024 [45]. Further, tight integration between GPUs and CPUs provide a large attack surface for GPU bugs [57].
Instead, Edera runs device drivers in a zone with the GPU attached to this zone, as shown in fig. 4. Applications run in their own zone and interact with this zone through an NVIDIA vGPU [46] that communicates with the GPU driver zone. The driver zone can be configured to passthrough the entire GPU to a single zone, or to share a GPU among multiple zones using partitioning. If any GPU user exploits a driver bug, only this GPU zone, and not the hypervisor or other user zones, will be compromised. Further, the zone can be configured to wipe GPU memory between different users of the GPU to prevent data leakage.
This same technique for driver isolation can be applied to other drivers. For example, network drivers managed by protect-network in fig. 3 can be moved into an Edera zone. Thus, network operations would be sent to this zone instead of the hypervisor. This moves these drivers out of the TCB and ensures that any vulnerability in the driver code can only be exploited within the zone and will not effect the hypervisor.
6 Analysis

### v2

@@ 2.4. Kubernetes
Kubernetes (maintainers, 2024e) is a system for automatically deploying, scaling, and managing containerized applications. It runs containers in pods that run on nodes. A container runtime manages the lifecycle of pods. Kubernetes can work with any container runtime that follows the Container Runtime Interface (CRI) specification (maintainers, 2024d). One such container runtime is runc, a lightweight runtime based on Docker. A container runtime interface, such as containerd provides higher-level control of containers on the node, including image management and container execution. The Kubernetes API server lets users query and change the state of Kubernetes across multiple nodes. A kubelet is an agent that runs on each node that registers the node with the API server and ensures that all pods expected by the API server are running.
@@
The Kubernetes ecosystem is composed of projects such as container runtimes, container image distribution services, and projects that add functionality or usability. The Cloud Native Computing Foundation alone hosts more than 200 open source projects in this ecosystem (Cloud Native Computing Foundation, 2025). The scope of this ecosystem means that any incompatible new technology will be at a disadvantage as it would lack compatibility with Kubernetes and related projects as well as pre-built OCI container images.
@@
Systems have been built to allow Kubernetes to manage VMs instead of or in addition to OS virtualized containers. This provides Kubernetes compatability to hypervisor virtualization. Many of these, including Infranetes (maintainers, 2016), rkt (maintainers, 2019a), and virtlet (maintainers, 2019b) require virtualization extensions. RunX (maintainers, 2021) uses a series of shell scripts to run VMs on Xen, but does not make optimizations in Xen or the VMM to improve performance. These projects demonstrate the desire for strongly isolated hypervisor virtualization systems compatible with Kubernetes, but none meet our performance and usability requirements.
3. Requirements
Our goal is to create a system with strong container isolation while ensuring compatibility with the container ecosystem and performance comparable to OS virtualization. To achieve this, our requirements are as follows.
@@ 3. Requirements
Eliminate the shared kernel. The system must support secure multi-tenancy, with the container acting as a security boundary. This means removing the kernel from the TCB so that a bug in the kernel is not sufficient to escape the container’s isolation.
@@
Use existing OCI images. Existing OCI container images must run without modification. Some OCI images have proprietary code which cannot be accessed by users to re-build the image. Even when the source is available, re-building images takes time and compute resources
Cloud compatibility. The system must run on commercially available systems without relying on specialized hardware like virtualization extensions or TEEs. Virtualization extensions are not available on all commercial cloud infrastructure as most cloud instances are already virtual machines, so access to virtualization extensions is only possible if the instance includes nested virtualization that passes through access to CPU features or if the instance provides bare metal access to a machine. Prices for a bare metal EC2 instance start at more than \$4/hour, compared to ¡ \$0.01 for the smallest non-bare metal instance.
@@
Memory Safety. The system should be written in a memory safe language to reduce the attack surface. Memory safety vastly reduces the number of memory leaks and memory corruption attacks.
Runtime performance. The system’s runtime performance including for workloads heavy in CPU, memory, or system call usage, must be similar to that in existing systems. This ensures that it will not require more resources or higher costs to run. The overhead of virtualization has been measured in previous work (Huber et al., 2011; Younge et al., 2011) with 5% being the lowest overhead reported, so we require that our performance be within 5% of running runc on Docker.
Startup performance. The system’s startup occurs just once per application, but speed is important when provisioning new containers and scaling the number of copies up and down. The startup for our system includes the VMM boot, guest OS boot and container runtime to be comparable with the container startup in Docker. Given this some overhead will be expected, but we set the goal of startup performance within 1 second of Docker, improving on the application start time in other hypervisor systems.
3.1. Threat model
@@
In multi-tenant environments, workloads from third parties are run on shared infrastructure. The infrastructure allows users to run their own code, but the operator does not want users to gain access to the underlying infrastructure or other workloads. Given this environment, for our threat model we assume an attacker can:
@@
  • Hardware attacks that require physical access to the machine.
@@
  • Side channel attacks, for instance those that use the temperature of the host machine. These attacks are unavoidable on shared infrastructure.
4. Design
@@
Edera uses a type-1 paravirtualized microkernel hypervisor that creates a root hardened zone that acts as a VMM, passing memory and CPU from the host machine to zones. The root hardened zone runs workloads in zones, which are fully functional VMs that run standard Linux for container workloads. The root hardened zone monitors for anomalies and can terminate other zones. The VMM controls and optimizes the memory and CPU usage of zones. Each zone is scheduled by the hypervisor with parameters from the root hardened zone like a user-space process within the hypervisor.
4.1. Hypervisor
The Edera hypervisor is composed of Xen (Barham et al., 2003) as the microkernel hypervisor, and our novel VMM. We call these elements collectively the Edera hypervisor. The Xen microkernel acts as the host kernel and can run directly on the host machine or as a guest OS in nested virtualization. It is a microkernel as the hypervisor only controls essential hardware interaction, while other components such as networking and device drivers are left for user-space processes in the root hardened zone and other isolated zones. Research has shown that microkernels have fewer bugs, as many bugs come from less used parts of the kernel (Li et al., 2017). The hypervisor is solely responsible for partitioning resources to each zone, and monitoring the usage of these resources.
Xen forms the base of Edera, but we utilize only the core components and not the orchestration or other tooling that are part of the Xen project. Xen uses MISRA C and boasts a mature codebase that makes it suitable for safety domains including automotive and aerospace (Xen Project, 2025a; Dorner Works, 2025; epam, 2025). Edera uses paravirtualization (PV) mode, with the hypervisor managing the CPU, MMU hardware page table, and network. There is also support for PV with hardware virtualization (PVH) mode, which takes advantage of virtualization hardware for performance improvements. As PVH mode requires virtualization hardware, examples in this paper use PV mode unless otherwise stated. Our build of Xen adds a small layer to simplify starting a new Xen domain, perform device enablement, and provide a Rust interface.
4.2. Root hardened zone
Edera’s novel VMM, which we call the root hardened zone, provides a system for VM orchestration, device management, dynamic resource allocation, zone initialization, and inter-process communication (IPC). The root hardened zone initializes the operating system in a zone, attaches disk and network devices to a zone, and performs RPC.
4.2.1. Resource manager
The Edera resource manager monitors resource usage, detects misbehavior, and performs dynamic resource allocation. The resource manager can also dynamically pin and unpin CPUs for each zone, and adjust the memory allocation. This dynamic configuration allows for more efficient resource allocation as zones can be assigned CPUs and memory pages as needed, rather than these resources remaining static after zone creation. This solves a common problem in containers, which need to be re-created if they run out of memory or CPU resources under static allocation.
Edera implements dynamic resource allocation that, to our knowledge, is the most expansive in any system. While it is inspired by memory ballooning (Chiang et al., 2013) that can add or remove memory based on usage, we take this a step further by providing prioritization for each zone. QubesOS performs similar prioritization, but its design assumes only one active VM at a time, and moves resources to this active VM. Edera provides resource limits and prioritization for memory pages and CPU cores for multiple concurrently running zones. These limits and priorities are initially set automatically, but can be updated by administrators as needed. Dynamic resource allocation takes advantage of Xen scheduling features to support non-standard memory designs such as non-uniform memory access (NUMA) (Xen Project, 2025b) to ensure efficient allocation even in environments with multiple CPUs.
4.2.2. Inter-process communication
IPC in Edera is handled with shared memory pages. When two zones need to communicate, for example when the root hardened zone needs to communicate with a newly created zone, the root hardened zone directs the hypervisor to allocate a page of memory and set it to be writable by both zones. The hypervisor can thus ensure that only these zones can access this page using the grant table, which is ceded to the hypervisor from the root hardened zone on initialization. To ensure the security of this communication, grant references can only be used by the zone they are given to. This process ensures that no zones, including the root hardened zone, can read communication between zones set up through IPC. Using the same process, driver zones attach to a zone by using the grant table to get shared pages with the zone. The device can then communicate to the attached zone, and no other zones, using this shared page.
4.3. Zones
@@
Each zone launched by the hypervisor runs a guest OS kernel. Using PV, the hypervisor runs in ring 0, and all guest OSs run in a less privileged ring (ring 1 for x86 and ring 3 on x86-64) (Nakajima et al., 2006). As most OSs expect to run in ring 0, the guest OS must be aware that it is running in a virtualized environment and support hypercalls. Most modern Linux distribution meet these requirements. These hypercalls (Barham et al., 2003) are used for privileged operations that need to run in ring 0 and pass along those privileged system calls to the hypervisor. This is supported by most OSs through a PV mode.
To optimize memory usage, guest kernels in Edera are placed in read-only memory that can be shared between multiple zones using the same operating system. These read-only kernel pages save system memory and prevent userspace programs from tampering with kernel functionality.
Edera supports running OCI images as the guest OS. Many OCI images are built as a minimal operating system interface for containers, and so provide a smaller footprint than installing a full Linux kernel. While not a true microkernel, these minimal images help improve the overhead of guest OSs in Edera.
4.3.1. Inter-domain messaging
@@
Edera uses inter-domain messaging (IDM) to send commands or processes to the zone (such as through the Kubernetes utility kubectl exec), in addition to sending logs and metadata back to the hypervisor. IDM uses a Xen byte channel (maintainers, 2024m) to send protobuf messages to other zones using IPC. Protobuf (shortened from protocol buffers) is a platform-neutral mechanism for serializing structured data. Its design includes protection against overloading the pipe by using bound checks and other limits to ensure safe operation.
4.3.2. Monitoring
@@
An init process in every zone is started when the zone is provisioned. This process runs in the background to monitor the VM. It sends all monitored events over the IDM channel to the root hardened zone to be logged and collected. If the init process is disabled, the root zone marks the zone as not responding and deprovisions it.
@@
5. Implementation
Figure 2. An overview of the services used by Edera to run a Kubernetes container.
@@ 5. Implementation
We implement Edera in Rust using the design from section 4 with built-in services for Kubernetes compatibility. By implementing it in Rust and utilizing MISRA C components from Xen, we avoid many memory safety bugs while achieving high performance. The implementation includes our build of the Xen hypervisor, the root hardened zone, and several services that allow use with Kubernetes for cluster orchestration. These services include a CRI compatible with the OCI runtime specification, daemons for zone monitoring, and a kube-apiserver for managing pods. For ease of use with Kubernetes, we create a custom guest OS kernel that is distributed as an OCI image. Thus, a user can specify this image in their Kubernetes configuration and automatically run their existing containers. The source for this implementation is available on request.
@@
We will walk through the system startup and running of a single Kubernetes pod in an Edera zone to illustrate the services used by the implementation, shown in fig. 2. First, the machine boots into the microkernel hypervisor. Xen microkernel starts the root hardened zone, which also acts as a Kubernetes host and runs protect-daemon. The hypervisor launches protect-orchestrator, a management component that ensures that all other daemons (protect-network, protect-daemon, and protect-cri) are running.
@@
protect-daemon implements the root hardened zone, keeping track of which zones are running each workload and managing drivers for the zones, while leaving resource management to the hypervisor. This is the daemon that creates, modifies, and removes zones. It manages an embedded key value store that tracks what is running in each zone where the zone is identified by a uuid that is assigned on zone creation. In addition, protect-daemon manages the drivers for each zone, including IDM. Using IDM, protect-daemon can pass commands from the API server (though protect-cri) to the init process in the zone. It can also receive messages from the zone, and pass these along to protect-cri.
@@
protect-orchestrator performs the initial system setup, and maintains running processes. It provides updates and health checks for running daemons including protect-daemon, protect-network, and protect-cri. In addition, it provides observability for all services running in the hypervisor, through tools like Prometheus (Authors, 2024).
@@
protect-network performs packet routing and monitors network activity. Zones have a basic ethernet interface, but cannot perform packet routing without going through protect-network. protect-network receives tasks from zones through protect-daemon, then performs them on behalf of the zone. It writes packets to the networking layer and can dial sockets for zones. Zones do not have direct network access, so protect-network sees all routed packets, ensuring that Edera can monitor all network activity.
@@
protect-cri provides a layer on top of any OCI-compliant CRI, and converts requests from the CRI to requests for protect-daemon. protect-cri reconciles the state between protect-daemon and the CRI to ensure that the zones running on protect-daemon match the Kubernetes pods expected by the CRI. It also transfers messages from zones through protect-daemon to other Kubernetes processes. protect-cri supports running some Kubernetes nodes in Edera zones and others on different infrastructure by coordinating with containerd to run the non-Edera nodes.
@@
Once the system is booted, a user may launch a pod from protect-cri. This triggers the kubelet (Kubernetes, 2024), which registers the pod with the kube-apiserver and launches a zone.
@@
Summary. In summary, Edera provides services to manage zones and communicate with Kubernetes. Each Kubernetes pod runs in its own Edera zone, ensuring no shared kernel state between pods. By routing all network traffic through protect-network and all IDM communication through protect-daemon, we ensure that zones can be monitored.
6. Analysis